Demography and Capital Flows:

Thomas F. Cooley Georgetown University | April 4, 2011

This version: April 19, 2011

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  1 / 1

Capital inflows: sign of success?

Larry Summmers, IMF, October 3, 2004:

There is a standard set of things that finance ministers of countries with significant current account deficits say. Perhaps the sharpest formulation is: “We live in a country that capital is trying to get into. Would you rather live in a country that capital is trying to get out of?”

Capital inflows: sign of success?

Larry Summmers, IMF, October 3, 2004:

There is a standard set of things that finance ministers of countries with significant current account deficits say. Perhaps the sharpest formulation is: “We live in a country that capital is trying to get into. Would you rather live in a country that capital is trying to get out of?”

Capital inflows: portent of peril?

Daniel Gross, New York Times, May 8, 2005:

[US] imbalances are eerily reminiscent of recent economic crises. Could we see a perfect storm [for the US economy]? If so, what would it look like?

Nouriel Roubini estimates that long-term interest rates in the US could rise by 200 basis points over a few months and the value of the dollar would fall.

Said Barry Eichengreen: “The result would not be a full-blown financial crisis most likely, but it would still be a major recession.”

Adds Jeffrey Frankel, “some of us have been warning of this hard-landing scenario for more than 20 years.”

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  3 / 1

Facts about capital flows

Facts about capital flows

What drives them?

▶ Business cycles ▶ Institutions ▶ Taxes and legal restrictions ▶ Commodity prices (oil) ▶ Political risk ▶ Exchange rates ▶ Anything that affects saving or investment

Some data to get us thinking...

Facts about capital flows

Facts: current accounts

Facts about capital flows

Facts: Net Foreign Assets